[Brahms-l] Shift report 20040322 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Mar 22 2004 - 00:14:35 EST
------------------------------------------------------------
Date: Mon Mar 22 00:14:34 2004, Shift: 16:00-24:00
Supervisor: BN, with: 
Log book#15 pages 94 - 94

Shift Summary:
------------------------------------------------------------
1. AGS had magnet problems, so store was kept for a very long time.  After the problem was fixed, beam came back.

2. Runs 11225-11233, made some silly mistakes in places, else more probably would have gotten done

3. No detector problems, things ran pretty smoothly.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Mar 22 00:14:50 2004

This archive was generated by hypermail 2.1.8 : Mon Mar 22 2004 - 00:15:12 EST